# RGB2Point: 3D Point Cloud Generation from Single RGB Images

### v3

@@ 1 Introduction
We analyze the impact of pre-trained weights on the effectiveness of the ViT for 3D point cloud reconstruction tasks. Moreover, we swap the image feature extractor part to a pre-trained ResNet50 [15] to evaluate the effect of ViT. Our method shows reconstruction improvements, including quantitative and qualitative results over existing works, as RGB2Point outperforms in Chamfer distance (39.26% and 51.15%) and Earth Mover’s distance (26.95% and 36.17%) compared to the current state-of-the-art. Also, our model shows 47.1% higher reconstruction quality in F-score compared to the diffusion-based method [33]. Contrary to previous work, our reconstructed point clouds show a 63.1% more consistently high quality for all object classes. We claim the following contributions: 1) providing a new model architecture that is VRAM efficient but also generates a high-quality point cloud from a single RGB image, 2) a Transformer model can generate higher quality 3D objects than a probabilistic denoising diffusion model. We will share the code and model upon acceptance.
@@ 5 Conclusions and Future Works
Based on the performance and efficiency of our model, it can be used as a prior or proxy before getting actual lidar-sensor scanned data. Scanning with lidar typically requires multi-view scans or object rotation. Our method generates a high-quality 3D point cloud from a single image in just 0.1 seconds, offering a fast and accurate alternative before full sensor-based scans, which require more time for post-processing. While multi-view approaches can improve performance, they increase memory use and runtime due to model complexity. In contrast, our method needs only 2.3GB of VRAM and 0.1 seconds on a standard GPU, providing an efficient tradeoff.
As a future work, RGB2Point has the potential to be adapted for generating specific domain objects. As we show in Sec. 4, combining our model with pre-trained weights on specific domains could be a solution for generating 3D point clouds from specific fields since RGB2Point can be trained on desktop-level hardware resources.
The scope of our research could be broadened to include the use of multi-view images, which could significantly enhance the accuracy of 3D point cloud generation. By incorporating cross-attention mechanisms among the different views, we can leverage the complementary information provided by multiple perspectives, thereby improving the overall fidelity and robustness of the generated 3D models.
Additionally, a potential improvement involves integrating a differentiable renderer into RGB2Point to apply RGB texture to the point cloud, enhancing visual fidelity. Given its VRAM efficiency, future work includes adapting this AR or VR device deployment model, enabling on-device usage. Furthermore, RGB2Point achieves point cloud generation in approximately 0.15 seconds per image. Considering this efficiency, the model could be quantized and optimized for edge devices, facilitating real-time single-image conditioned generation. Such capabilities could benefit applied robotics, particularly path planning and object evasion maneuvers.

### v4

@@ 1 Introduction
We analyze the impact of pre-trained weights on the effectiveness of the ViT for 3D point cloud reconstruction tasks. Moreover, we swap the image feature extractor part to a pre-trained ResNet50 [14] to evaluate the effect of ViT. Our method shows reconstruction improvements, including quantitative and qualitative results over existing works, as RGB2Point outperforms in Chamfer distance (39.26% and 51.15%) and Earth Mover’s distance (26.95% and 36.17%) compared to the current state-of-the-art. Also, our model shows 47.1% higher reconstruction quality in F-score compared to the diffusion-based method [32]. Contrary to previous work, our reconstructed point clouds show a 63.1% more consistently high quality for all object classes. We claim the following contributions:
  • We propose a new model architecture that is VRAM efficient but also generates a high-quality point cloud from a single RGB image.
  • We show Transformer model can generate higher quality 3D objects than a probabilistic denoising diffusion model.
@@ 5 Conclusions and Future Works
Based on the performance and efficiency of our model, it can be used as a prior before getting actual lidar-sensor scanning, which requires multi-view scans. Our method generates a high-quality 3D point cloud from a single image in just 0.1 seconds, offering a fast and accurate alternative.
Future work could adapt RGB2Point for generating domain-specific objects by combining it with pre-trained weights, enabling 3D point cloud generation on desktop-level hardware. Expanding to multi-view images, with cross-attention mechanisms, could improve accuracy by leveraging complementary information from different perspectives, enhancing fidelity and robustness. Additionally, integrating a differentiable renderer for RGB texturing would increase visual quality. With VRAM efficiency, RGB2Point could be optimized for AR/VR deployment, achieving real-time, on-device use. Its fast generation time (0.15 seconds per image) makes it ideal for robotics tasks like path planning and object evasion.
@@
Acknowledgement
This work was supported in part by the U.S. National Science Foundation under awards No. 2412928 and 2417510. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work is based upon efforts supported by the USDA-NIFA grants 2024-67021-42879 “Improving Forest Management Through Automated Measurement of Tree Geometry” and 2023-08563 “Optimizing soybean canopy architecture for efficient light and water use under climate change”. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government or NRCS. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.